Sarah had a test on Thursday. She studied for the test for 2/3 of an hour on Monday, 3/4 of an hour on Tuesday, and 1 1/3 of an hour on Wednesday. How many hours did she study for the science test over these three days
Answer by rvfer(117)

2/3+3/4+4/3=
8/12+9/12+16/12=33/12=2 3/4 hrs